a.     Suppose the number of employed people in Pakistan is 20.5 million. The unemployment rate in this economy is 8.8 percent, and the labor force participation rate is 65 percent. (1.5 Marks)

i.       What is the size of the labor force and working-age population?

ii.     How many people are unemployed? 



1
Expert's answer
2020-11-30T16:41:05-0500

i) employment rate "=100\%-8.8\%=91.2\%"

total labor force "=20.5\times \frac{100}{91.2}=22.47million"

total adult population"=22.47\times \frac{100}{65} = 34.58million"

ii) number of unemployed "=8.8\%" of "22.47million = 1.97million"